# SLIM: Let LLM Learn More and Forget Less with Soft LoRA and Identity Mixture

### v1

@@ Abstract
Although many efforts have been made, it is still a challenge to balance the training budget, downstream performance, and the general capabilities of the LLMs in many applications. Training the whole model for downstream tasks is expensive, and could easily result in catastrophic forgetting. By introducing parameter-efficient fine-tuning (PEFT), the training cost could be reduced, but it still suffers from forgetting, and limits the learning on the downstream tasks. To efficiently fine-tune the LLMs with less limitation to their downstream performance while mitigating the forgetting of general capabilities, we propose a novel mixture of expert (MoE) framework based on Soft LoRA and Identity Mixture (SLIM), that allows dynamic routing between LoRA adapters and skipping connection, enables the suppression of forgetting. We adopt weight-yielding with sliding clustering for better out-of-domain distinguish to enhance the routing. We also propose to convert the mixture of low-rank adapters to the model merging formulation and introduce fast dynamic merging of LoRA adapters to keep the general capabilities of the base model. Extensive experiments demonstrate that the proposed SLIM is comparable to the state-of-the-art PEFT approaches on the downstream tasks while achieving the leading performance in mitigating catastrophic forgetting.
SLIM: Let LLM Learn More and Forget Less with Soft LoRA and Identity Mixture
Jiayi Han1,2, Liang Du3, Hongwei Du1,2, Xiangguo Zhou1,2, Yiwen Wu1,2, Weibo Zheng1,2, Donghong Han4, 1Inspur Genresoft Co. Ltd., Inspur Group Co. Ltd. 2Shandong Key Laboratory of Automated Complex Network Software Construction 3Interactive Entertainment Group, Tencent Inc. 4Northeastern University, China Correspondence: hanjiayi@inspur.com, handonghong@cse.neu.edu.cn
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated extraordinary general capabilities, and are widely used for question-answering, code generation, mathematics, and other fields. However, the increasing utilization of LLMs in specialized domains needs to support tasks with a mixture of world and specific knowledge. Meanwhile, since the scale of the LLMs is large, fine-tuning the whole model could be expensive, and may not be affordable for many users. Furthermore, after fine-tuning on the downstream tasks, the LLMs may lose the general capability, due to catastrophic forgetting. How to reduce the training cost while mitigating catastrophic forgetting without loss of learning capacity, is an essential problem in the utilization of LLMs, especially when utilized in a scenario where both common and professional capabilities are needed.
@@
An intuitive solution is to train only a small bunch of parameters. For example, low-rank adaptation (LoRA) Hu et al. (2022), a parameter-efficient fine-tuning (PEFT) approach, is validated to forget fewer pieces of knowledge compared with fine-tuning the full weight of the base model, with limitations of learning capacity Biderman et al. (2024). Chaudhry et al. (2020) propose to optimize the model in the orthogonal subspace with low-rank adaptation to alleviate forgetting. Singh et al. (2020) propose to insert calibrating modules to the base model for learning the downstream tasks. Liang et al. (2023) utilize learnable mask prompts to fine-tune the CLIP without loss of text-vision alignment. These approaches attempt to introduce minimal changes to the base model to keep its raw capacity. Some approaches utilize data re-play to reduce forgetting. For example, Glorioso et al. (2024) utilize a 60% data replay for continual training. Zhu et al. (2024) utilize a 30% re-play rate in the continual pre-training for the coding model. Despite the fact that data replay could effectively reduce the influence of forgetting, the large amount of training data is still a burden, and the pre-training data may not be accessible for downstream fine-tuning.
In this work, we propose a novel MoE architecture with Soft LoRA and Identity Mixture (SLIM), a mixture of LoRA adapters and identity layers. By introducing identity layers as experts, the model has the potential of routing the samples out of the downstream distribution to the identity layers, thus avoiding the influence of downstream tasks. To encourage the model to dynamically route the samples from the downstream tasks to the LoRA adapters, while assigning the samples out of the downstream distribution to the identity layers to avoid forgetting, we propose weight yielding with sliding clustering to estimate the distribution of the samples from the downstream tasks and modify the routing weight according to the consistency between the input sample and the distribution. Inspired by DARE Yu et al. (2024), we further convert the low-rank adaptation to model merging task and propose a fast merging approach that dynamically merges the mixture of LoRA adapters to the base model. We conduct experiments on various downstream datasets while validating the model performance on multiple general tasks. Extensive experiments show that the proposed SLIM could achieve SOTA PEFT performance, and significantly mitigate catastrophic forgetting, as demonstrated in Fig. 1. Our main contributions could be summarized as follows:
@@
    We propose weight yielding with sliding clustering that enhances the dynamic routing between the identity layers and the LoRA adapters for better downstream performance and catastrophic forgetting suppression.
@@
    We propose dynamic merging that converts the MoE low-rank adaptation to model merging, which effectively eliminates catastrophic forgetting without any data replay.
@@ 2.1 Parameter-efficient fine-tuning (PEFT)
Adapter modules introduce additional adaptive layers to the pre-trained model for fine-tuning. Serial Adapter Houlsby et al. (2019) adding adapter modules cascaded to the self-attention layers and the FFN layers. AdapterFusion Pfeiffer et al. (2021) further proposed to insert the adapters only after the normalization layers for the FFN layers to further boost training efficiency. Some approaches also introduce parallel adapters to the transformer layers He et al. (2021); Lei et al. (2023).
@@
LoRA-styled adapters assume that the change of the full-weight matrix could be approximated by a low-rank matrix Hu et al. (2022). AdaLoRA Zhang et al. (2024) proposes to adaptively estimate the parameters’ importance and modify the adapters’ rank. DoRA Liu et al. (2024) further decomposes the pre-trained weights to scale and direction, proposing only to update the direction of the pre-trained weights. MoRA Jiang et al. (2024b) proposes to increase the rank of the adapters with the same number of parameters. Inspired by the Mixtral 8$\times$7B model Jiang et al. (2024a), LoRAMoE Dou et al. (2024) and MixLoRA Li et al. (2024) propose to mixing the LoRA adapters with dynamic routing. LoRAMoE further introduces a contrastive loss function to mitigate knowledge forgetting during downstream training.
@@ 2.2 Task incremental learning
Replay-based methods, such as Experience Replay (ER) and Generative Replay, mitigate forgetting by reintroducing samples from previous tasks during training on new tasks. For example, Rolnick et al. (2019) leverages behavior cloning and off-policy learning from the replay for training stability. Chaudhry et al. (2019) maintain a memory bank during training, and sample from the bank when training the consequent batches. Many approaches further modify the hyperparameters and training schema (e.g. learning rate, re-warm up) along with the data re-play during the continual pre-training Zhu et al. (2024); Glorioso et al. (2024); Ibrahim et al. (2024).
@@
Some approaches adopt extra regularization terms to the adapted weights to mitigate forgetting. For example, Chaudhry et al. (2020) proposes to fine-tune the base model on the orthogonal sub-space of its parameters to avoid the influence of the tasks trained previously. LoRAMoE Dou et al. (2024) proposes to leverage contrastive loss during fine-tuning to alleviate the loss of world knowledge. Wang et al. (2021) introduce Adam-NSCL that forces the update of the network parameters to lie in the null space of the previous tasks for capability reservation.
@@ 3 Method
Figure 2: The overall framework of the proposed approach. “Iden” represents the identity layer. The input series are first sent to the router to predict the routing weight of each identity layer and LoRA adapters. We also calculate the distance from the input to the clusters, and the routing weight is modified with the distance by the weight yielder. Then we select top-$K$ experts according to the yielded weight. The final output is achieved by the dynamic merging that replaces the normal weighted sum of the MoE methods.
@@ 3.1 Overview of LoRA-styled adapters
$\begin{array}[]{l}y=Wx+b+\text{LoRA}(x)=Wx+b+BAx,\end{array}$ (1)
@@
Although LoRA is efficient and does not directly modify the weight of the base model, it still influences the output and therefore impacts the generalization of the LLMs. Since LoRA adapters are independent modules of the base model, a direct solution for recovering the generalization is to discard the LoRA adapters from the base model when not needed. We formulate this alternative implementation of LoRA adapters as in Eq. 2:
@@
in which $\hat{R}:\mathbb{R}^{C}\rightarrow\{0,1\}$ is an expert (e.g., human) routing operation that determines if the LoRA adapter should be discarded. Since it is not feasible to let users decide whether to utilize the LoRA adapter for each individual instruction, we propose a dynamic approach as illustrated in the following.
@@ 3.2 Mixture of experts as soft $\hat{R}$
Similar to the MoE architecture of the Mixtral models, the mixture of LoRA adapters also introduces a routing layer to assign the input tokens to multiple LoRA adapters. Formally, given input $x\in\mathbb{R}^{B,N,C_{in}}$ in which $B,N,C_{in}$ represent the batch size, length of input tokens, and the number of channels, a routing layer $R$ is adopted to assign the tokens to the LoRA experts as in Eq. 3:
$\hat{w}_{i}=\left\{\begin{array}[]{l}\text{Softmax}\left(R(x)\right)_{i},\mathcal{R}(R(x)_{i})\leq 2\\ 0,otherwise\end{array}\right.$ (3)
@@
$\left\{\begin{array}[]{l}y=Wx+\frac{1}{Z}\sum\limits_{i}\hat{w}_{i}\text{LoRA}_{i}(x)+b\\ Z=\sum_{i}\textbf{1}(\mathcal{R}(\hat{w}_{i})\leq 2)w_{i}\end{array}.\right.$ (4)
To automatically assign the tokens to the adapters, we replace the expert routing operation $\hat{R}$ with a learnable routing layer conditioned on the input as:
@@
in which $f_{i}$ could either be a LoRA adapter or an identity layer that satisfies $f(x)=x$, and the routing weight $\hat{R}(x)$ satisfies $\sum_{i}\hat{R}(x)_{i}=1$. Note that we do not utilize the 0 matrix in our method. By setting $\hat{R}(x)_{i}=\hat{w}_{i}/Z$, Eq. 6 could be transformed to the same formulation with Eq. 4.
@@ 3.3 Weight yielding with sliding clustering
To identify the samples occluded from the target domain, we suppose the distribution of the input samples is a mixture of Gaussian distribution. We calculate the distance from the input sample to the nearest cluster. We suppose that if the distance is large, the sample is rarely probable to be sampled from the target domain, therefore it should be processed by the identity expert. During training, we assign each input sample to a cluster and update its center. We also estimate the variance of each channel of the cluster. We do not update the clusters during inference.
Formally, we randomly initialize $N$ cluster centers $\mathcal{C}=\{c_{i}\in\mathbb{R}^{C}\}$. For each input sample $x\in\mathbb{R}^{N,C}$, we assign the sample to the nearest cluster as follows:
$\begin{array}[]{l}\text{idx}|_{x}=\mathop{\text{argmin}}\limits_{i}\left(||c_{i}-\frac{1}{N}\sum\limits_{k}x_{k,:}||_{2}\right).\end{array}$ (7)
We then yield the routing weights as follows:
@@
in which $\text{TYPE}(\cdot)$ return the type of the layer, $\tilde{R}:\mathbb{R}^{C}\rightarrow\mathbb{R}$ calculate the routing logits of the input, and the distance $d$ could be calculated as follows:
@@
The routing weight could finally be formulated as follows:
@@ 3.4 Mixing LoRA adapters with dynamic merging
DARE Yu et al. (2024) proposes to merge the LLMs with the same pre-trained parameters to obtain the capacity of the merged LLMs. DARE supposes the weights of the fine-tuned models could be formulated as $\theta_{\text{SFT}}=\theta_{\text{PRE}}+\tau$, in which $\theta_{\text{SFT}},\theta_{\text{PRE}}$ represent the weight of the fine-tuned model and the pre-trained model. $\tau$ is named as “task vector”. As depicted in DARE, LLM have the following natures:
@@
Inspired by the findings of DARE, we propose to convert the mixing of multiple LoRA adapters to dynamic model merging to further recover the generalization of the LLM. DARE randomly sets the elements of $\tau$ to 00 to formulate $\theta_{\text{DARE}}$. The merging of the model could be formulated as follows:
$\theta_{M}&=\theta_{\text{PRE}}+\lambda\sum\limits_{k}(\theta_{\text{DARE}}^{t_{k}}-\theta_{\text{PRE}})\\ &=\theta_{\text{PRE}}+\frac{\lambda}{1-p}\sum\limits_{k}M_{t_{k}}(\theta^{t_{k}}_{\text{SFT}}-\theta_{\text{PRE}})\end{aligned},$ (11)
in which $\otimes$ represents the Hadamard product, and $p\in(0,1)$ is the masking rate. We formulate the merging of the general instruction tuning model and the downstream model trained by LoRA as:
@@
in which $\theta_{M},\theta_{\text{PRE}},\theta_{\text{Ins}}$ represent the weight of the merged model, the pre-trained LLM, and the instruction tuning model, $M_{L}$ and $M_{A}$ are binary masks. To simplify the merging procedure, we need to eliminate the $\theta_{\text{PRE}}$ from the equation to avoid access to the pre-training model during inference. Suppose $M_{L}$ and $M_{A}$ satisfy $XOR(M_{L},M_{A})=\textbf{1}$. Setting $\lambda=1-p$, Eq. 12 could be formulated as follows:
$\theta_{M}&=\theta_{\text{PRE}}+\frac{\lambda}{1-p}M_{L}\otimes(\theta_{\text{Ins}}-\theta_{\text{PRE}})\\ &+\frac{\lambda}{1-p}M_{A}\otimes(\theta_{\text{Ins}}+BA-\theta_{\text{PRE}})\\ &=\theta_{\text{Ins}}+M_{A}\otimes(BA)\end{aligned}.$ (13)
@@
$\begin{array}[]{l}\theta_{M}=\theta_{Ins}+\sum_{i}\hat{w}_{i}M_{A}^{i}\otimes(B_{i}A_{i})\end{array}.$ (14)
It is worth noting that we ignore the identity layers for simplicity. Although Eq. 14 could mimic the DARE approach to the mixture of LoRA experts, calculation of the full $BA$ for all LoRA adapters is required, resulting in multiple $d_{1}\times d_{2}$ matrices, which is not computationally efficient. To address this issue, we propose a fast implementation to approximate the random masking.
Formally, denote the row set of the masked elements as $\mathcal{S}_{\text{row}}=\{i|\exists j\in[0,d_{2}],M_{i,j}=1\}$ and the set of the column set $\mathcal{S}_{\text{col}}=\{j|i\in[0,d_{1}],M_{i,j}=1\}$, we could approximate the masking as follows:
@@
in which $\text{MASK}(a,b)$ is a function that masks the rows of $a$ according to the index set $b$, which could be formulated as follows:
$\text{MASK}(a,b)_{i,:}\left\{\begin{array}[]{l}\textbf{0},i\in b\\ a_{i,:},otherwise\end{array}.\right.$ (16)
Compared with masking with Hadamard product, this approach will set the whole column (row) to 0, which dramatically increases the masking ratio. To address this, we randomly sample a subset of $\mathcal{S}_{\text{row}}$ and $\mathcal{S}_{\text{col}}$ to fit the original masking ratio. It is worth noting that during inference, we directly masking the rows and columns, instead of the approximation approach mentioned above.
We also introduce $L_{1}$ regularization to the adapters. Denote $l$ and $k$ as the index of layers and LoRA adapters, it could be formulated as:

### v2

@@ Abstract
Despite the recent efforts from the NLP community, balancing the training budget, downstream performance, and general capabilities of large language models (LLM) remains a challenge in many applications. Training the entire model for downstream tasks is expensive, and could easily result in catastrophic forgetting. Parameter-efficient fine-tuning (PEFT) could reduce the training cost, but it still suffers from forgetting, and limits the learning on the downstream tasks. To address the aforementioned issues, we propose a novel mixture of experts (MoE) framework based on Soft LoRA and Identity Mixture (SLIM). SLIM allows dynamic routing between LoRA adapters and identity layers, thus enabling the bypass of LoRA adapters to suppress forgetting of general capacity. We adopt weight yielding with sliding clustering for better out-of-domain distinguish to enhance the routing. We also convert the mixture of LoRA adapters to the model merging formulation and introduce dynamic merging with its fast implementation for LoRA adapters to keep the general capabilities. Extensive experiments demonstrate that the proposed SLIM is comparable to the state-of-the-art PEFT approaches on the downstream tasks while achieving the leading performance in mitigating catastrophic forgetting. We plan to open-source the code upon publication.
SLIM: Let LLM Learn More and Forget Less with
Soft LoRA and Identity Mixture
Jiayi Han1,2, Liang Du3, Hongwei Du1,2, Xiangguo Zhou1,2, Yiwen Wu1,2, Weibo Zheng1,2, Donghong Han4, 1Inspur Genersoft Co. Ltd., Inspur Group Co. Ltd. 2Shandong Key Laboratory of Automated Complex Network Software Construction 3Interactive Entertainment Group, Tencent Inc. 4Northeastern University, China Correspondence: hanjiayi@inspur.com, handonghong@cse.neu.edu.cn
@@ 1 Introduction
In recent years, large language models (LLM) have demonstrated extraordinary general capabilities, and are widely used for question-answering, code generation, mathematics, and other fields. However, the increasing utilization of LLM in specialized domains needs to support tasks with a mixture of world and specific knowledge. Meanwhile, since the scale of the LLM is large, fine-tuning the entire model could be expensive, and thus may not be affordable for many users. Furthermore, after fine-tuning the downstream tasks, the LLM may lose the general capability, due to catastrophic forgetting. How to reduce the training cost while mitigating catastrophic forgetting without loss of learning capacity, is an essential problem in the utilization of LLM, especially when utilized in a scenario where both common and professional capabilities are needed.
@@
An intuitive solution is to train only a small bunch of parameters. For example, low-rank adaptation (LoRA) Hu et al. (2022), a widely used PEFT approach, is validated to forget fewer pieces of knowledge compared with fine-tuning the full weight of the base model, with limitations of learning capacity Biderman et al. (2024). Chaudhry et al. (2020) propose to optimize the model in the orthogonal subspace with low-rank adaptation to alleviate forgetting. Singh et al. (2020) propose to insert calibrating modules to the base model for learning the downstream tasks. Liang et al. (2023) utilize learnable mask prompts to fine-tune the CLIP without loss of text-vision alignment. These approaches attempt to introduce minimal changes to the base model to keep its raw capacity. Some approaches utilize data replay to reduce forgetting. For example, Glorioso et al. (2024) utilize a 60% data replay for continual training. Zhu et al. (2024) utilize a 30% replay rate in the continual pre-training for the coding model. Despite the fact that data replay could effectively reduce the influence of forgetting, the large amount of training data is still a burden, and the pre-training data may not be accessible for downstream fine-tuning.
In this work, we propose a novel MoE architecture with Soft LoRA and Identity Mixture (SLIM), a mixture of LoRA adapters and identity layers. By introducing identity layers as experts, the model is enabled to activate these layers as information “highways” to skip the LoRA adapters, thus avoiding the influence of downstream tasks. We further propose weight yielding with sliding clustering to estimate the distribution of the samples from the downstream tasks and modify the routing weight according to the consistency between the input sample and the distribution. This mechanism encourages the model to dynamically route the samples from the downstream tasks to the LoRA adapters while assigning the samples out of the downstream distribution to the identity layers to alleviate forgetting. Inspired by DARE Yu et al. (2024), we convert the low-rank adaptation to model merging formulation and propose a dynamic merging approach and its fast implementation to mix LoRA adapters to the base model for forgetting alleviation. We conduct experiments on various downstream datasets while validating the model performance on multiple general tasks. Extensive experiments show that the proposed SLIM could achieve SOTA PEFT performance, and significantly mitigate catastrophic forgetting, as demonstrated in Fig. 1. Our main contributions could be summarized as follows:
@@
    We propose a weight-yielding mechanism with sliding clustering to enhance dynamic routing between identity layers and LoRA adapters, improving input distribution distinction and enhancing general competence preservation.
@@
    We propose dynamic merging that converts the MoE low-rank adaptation to model merging, which effectively mitigates catastrophic forgetting without any data replay.
@@ 2.1 Parameter-efficient fine-tuning (PEFT)
Adapter modules introduce additional adaptive layers to the pre-trained model for fine-tuning. Serial Adapter Houlsby et al. (2019) introduces adapter modules cascaded to the self-attention layers and the FFN layers. AdapterFusion Pfeiffer et al. (2021) further proposes to insert the adapters only after the normalization layers for the FFN layers to further boost training efficiency. Some approaches also introduce parallel adapters to the transformer layers He et al. (2021); Lei et al. (2023).
@@
LoRA-styled adapters assume that the change of the full-weight matrix could be approximated by a low-rank matrix Hu et al. (2022). AdaLoRA Zhang et al. (2024) proposes to adaptively estimate the parameters’ importance and modify the adapters’ rank. DoRA Liu et al. (2024) further decomposes the pre-trained weights to scale and direction, proposing only to update the direction of the pre-trained weights. MoRA Jiang et al. (2024b) proposes to increase the rank of the adapters with the same number of parameters. Inspired by the Mixtral 8$\times$7B model Jiang et al. (2024a), LoRAMoE Dou et al. (2024) and MixLoRA Li et al. (2024) propose mixing the LoRA adapters with dynamic routing. LoRAMoE further introduces a contrastive loss function to mitigate knowledge forgetting during downstream fine-tuning.
@@ 2.2 Task incremental learning
Replay-based methods, such as Experience Replay (ER) and Generative Replay, mitigate forgetting by reintroducing samples from previous tasks during training on new tasks. For example, Rolnick et al. (2019) leverage behaviour cloning and off-policy learning from the replay for training stability. Chaudhry et al. (2019) maintain a memory bank during training, and sample from the bank when training the consequent batches. Many approaches further modify the hyperparameters and training schema (e.g. learning rate, re-warm up) along with the data replay during the continual pre-training Zhu et al. (2024); Glorioso et al. (2024); Ibrahim et al. (2024).
@@
Some approaches adopt extra regularization terms to the adapted weights to mitigate forgetting. For example, Chaudhry et al. (2020) propose to fine-tune the base model on the orthogonal sub-space of its parameters to avoid the influence of the tasks trained previously. Wang et al. (2021) introduce Adam-NSCL that forces the update of the network parameters to lie in the null space of the previous tasks for capability reservation. Zhai et al. (2024) propose a selective patch-level distillation approach to adaptively balance plasticity and stability.
@@ 3 Method
Figure 2: The overall framework of the proposed approach. “Iden” denotes the identity layer, which could be activated by the router as an information “highway” to bypass the LoRA adapters (see Sec. 3.2 for details), thereby preserving general capabilities. First, the router predicts the routing weight for each expert. Second, we calculate the distance from the input to the clusters and yield the routing weight accordingly. Then we activate the top-$K$ experts according to the yielded routing weight and fuse their output via dynamic merging. We remove the connections from the input to the weight matrices ($W$, LoRA adapters) and the “Iden” in the figure for simplicity.
@@ 3.1 Overview of LoRA-styled adapters
$\begin{array}[]{l}y=Wx+b+\mathcal{E}(x)=Wx+b+BAx,\end{array}$ (1)
@@
Although LoRA is efficient and does not directly modify the weight of the base model, it still influences the output and therefore impacts the generalization of the LLM. Since LoRA adapters are additional modules to the base model, an intuitive solution for recovering the generalization is to discard the LoRA adapters from the fine-tuned model when not needed. We formulate this alternative implementation of LoRA adapters as in Eq. 2:
@@
in which $\hat{R}\in\{0,1\}$ is an expert (e.g., human) routing operation that determines if the LoRA adapter should be discarded. Since it is not feasible to let users decide whether to utilize the LoRA adapter for each individual instruction, we propose a dynamic approach as illustrated in the following.
@@ 3.2 Mixture of experts as soft $\hat{R}$
Similar to the MoE architecture of the Mixtral models, the mixture of LoRA adapters also introduces a routing layer to assign the input tokens to multiple LoRA adapters. Formally, given $M$ LoRA adapters and input token $x\in\mathbb{R}^{C}$ in which $C$ represent the number of channels, a routing layer $R:\mathbb{R}^{C}\rightarrow\mathbb{R}^{M}$ is adopted to assign the token to the LoRA experts as in Eq. 3:
$\hat{w}_{i}=\left\{\begin{array}[]{l}\text{Softmax}\left(R(x)\right)_{i},\mathcal{R}(R(x)_{i})\leq K\\ 0,otherwise\end{array}\right.$ (3)
@@
$\left\{\begin{array}[]{l}y=Wx+\frac{1}{Z}\sum\limits_{i}\hat{w}_{i}\mathcal{E}_{i}(x)+b\\ Z=\sum_{i}\hat{w}_{i}\end{array}.\right.$ (4)
To dynamically assign the token to the adapters, we implement the expert routing operation $\hat{R}$ with a learnable routing layer, formulated as follows:
@@
in which $f_{i}$ could either be a LoRA adapter, or an identity layer $\mathcal{F}$ that satisfies
$g(x)+\mathcal{F}(x)=g(x),$ (7)
which holds for any function $g$ and any input $x$. By setting $\hat{R}(x)_{i}=\hat{w}_{i}/Z$, and considering both LoRA adapters and identity layers as experts $\{\mathcal{E}_{k}\}$, Eq. 6 could be transformed to the same formulation with Eq. 4.
@@ 3.3 Weight yielding with sliding clustering
To identify the samples occluded from the downstream tasks, we assume the distribution of the input samples during downstream fine-tuning is a mixture of Gaussian distribution. Therefore, we calculate the distance from the input sample to the nearest cluster. We suppose that if the distance is large, the sample is rarely probable to be sampled from the downstream distribution, therefore it should be processed by the identity layer. In the training phase, we assign each input sample to a cluster and update the cluster center accordingly. We also estimate the variance of each channel of the cluster. We do not update the clusters during inference.
Formally, we randomly initialize a set of cluster centers $\mathcal{C}=\{c_{i}\in\mathbb{R}^{C}\}$. For each input sample $x\in\mathbb{R}^{C,N}$, we assign the sample to the nearest cluster as follows:
$\begin{array}[]{l}\text{idx}|_{x}=\mathop{\text{argmin}}\limits_{i}\Big{|}c_{i}-\frac{1}{N}\sum\limits_{k}x_{:,k}\Big{|}.\end{array}$ (8)
We then yield the routing weights as:
@@
in which $\text{TYPE}(\cdot)$ return the type of the expert, $\tilde{R}:\mathbb{R}^{C}\rightarrow\mathbb{R}^{M}$ calculate the routing logits of the input, and the distance $d$ could be calculated as follows:
@@
The yielded routing weight could finally be formulated as follows:
@@ 3.4 Mixing LoRA adapters with dynamic merging
DARE Yu et al. (2024) proposes to merge multiple LLMs with the same pre-trained parameters to obtain the capacity of the merged LLMs. DARE supposes the weights of the fine-tuned models could be formulated as $\theta_{\text{SFT}}=\theta_{\text{PRE}}+\tau$, in which $\theta_{\text{SFT}},\theta_{\text{PRE}}$ represent the weight of the fine-tuned model and the pre-trained model. $\tau$ is named as “task vector”. As depicted in DARE, LLM has the following natures:
@@
Inspired by the findings of DARE, we propose to convert the mixing of multiple LoRA adapters to dynamic model merging to further recover the generalization of the LLM. DARE randomly sets the elements of $\tau$ to 00 to formulate $\theta_{\text{DARE}}$. The merging of the models could be formulated as follows:
$\theta_{M}&=\theta_{\text{PRE}}+\lambda\sum\limits_{k}(\theta_{\text{DARE}}^{t_{k}}-\theta_{\text{PRE}})\\ &=\theta_{\text{PRE}}+\frac{\lambda}{1-p}\sum\limits_{k}M_{t_{k}}\otimes(\theta^{t_{k}}_{\text{SFT}}-\theta_{\text{PRE}})\end{aligned},$ (12)
in which $\otimes$ represents the Hadamard product, $p\in(0,1)$ is the masking rate, $M_{t_{k}}$ is the binary mask for task $k$. $\theta_{M},\theta_{\text{PRE}},\theta_{\text{SFT}}^{t_{k}}$ represent the weight of the merged model, the pre-trained model, and the fine-tuned model for task $k$. We formulate the merging of the general instruction tuning model and the downstream fine-tuned model as:
@@
in which $M_{L}$ and $M_{A}$ are random binary masks, and $\theta_{\text{Ins}}$ is the instruction-tuning model. We need to eliminate the $\theta_{\text{PRE}}$ from the equation to avoid access to the pre-training model during inference. Suppose $M_{L}$ and $M_{A}$ satisfy $XOR(M_{L},M_{A})=\textbf{1}$. Setting $\lambda=1-p$, Eq. 13 could be simplified as follows:
$\theta_{M}$ $=\theta_{\text{PRE}}+\frac{\lambda}{1-p}M_{L}\otimes(\theta_{\text{Ins}}-\theta_{\text{PRE}})$ (14)
$\quad+\frac{\lambda}{1-p}M_{A}\otimes(\theta_{\text{Ins}}+BA-\theta_{\text{PRE}})$
$=\theta_{\text{Ins}}+M_{A}\otimes(BA).$
@@
$\begin{array}[]{l}\theta_{M}=\theta_{Ins}+\frac{1}{Z}\sum_{i}\hat{w}_{i}M_{A}^{i}\otimes(B_{i}A_{i})\end{array}.$ (15)
For simplicity, we ignore the identity layers in this equation. Although Eq. 15 could mimic the DARE approach to the mixture of LoRA experts, the calculation of the full-size $BA$ for all LoRA adapters is required, resulting in multiple $d_{1}\times d_{2}$ matrices, which is not computationally efficient. To address this issue, we propose a fast implementation to approximate the random masking. Formally, denote the row set of the masked elements as $\mathcal{S}_{\text{row}}=\{i|\exists j\in[0,d_{2}],M_{i,j}=0\}$ and the column set $\mathcal{S}_{\text{col}}=\{j|\exists i\in[0,d_{1}],M_{i,j}=0\}$, we could approximate the masking as follows:
@@
[…]
$\text{MASK}(\textbf{a},\textbf{b})_{i,:}\left\{\begin{array}[]{l}\textbf{0},i\in\textbf{b}\\ \textbf{a}_{i,:},otherwise\end{array}.\right.$ (17)
Compared with masking with Hadamard product, this approach will set the entire row (column) to 0, which dramatically increases the masking ratio. To address this, we randomly sample a subset of $\mathcal{S}_{\text{row}}$ and $\mathcal{S}_{\text{col}}$ to fit the original masking ratio. It is worth noting that, for simplicity of implementation, we directly mask the rows and columns instead of using the approximation approach mentioned above.
We introduce $L_{1}$ regularization to the adapters. Denote $l$ and $k$ as the index of layers and the attached LoRA adapters, it could be formulated as: